OpenStudy (anonymous):

If it requires 23.4 milliliters of 0.65 molar barium hydroxide to neutralize 42.5 milliliters of nitric acid, solve for the molarity of the nitric acid solution. Show all of the work used to solve this problem. Unbalanced equation: Ba(OH)2 + HNO3 Ba(NO3)2 + H2O

OpenStudy (anonymous):

Hi there! The balanced equation is: Ba(OH)2 + 2 HNO3 -----> Ba(NO3)2 + 2H2O How you have to think this problem: look at the balanced equation; you will observe that 1 mole of Ba(OH)2 is reacting with 2 moles of HNO3. You analyse the datas that are given in the problem. You have the volume of the solution of Ba(OH)2 used in the reaction and its molar concentration. You can easily calcule the moles of Ba(OH)2 found in the solution. cM=0.65M; V=23.4 mL solution of Ba(OH)2 n=cM*V=0.65*23.4=15.21mmoles Ba(OH)2 You know the volume of HNO3 that was neutralised by Ba(OH)2. You are required to calculate the molar concentration of the HNO3 solution, but you do not have the moles of HNO3, right? You can calculate the moles of HNO3 is you know the moles of Ba(OH)2. n'=2*n=2*15.21=30.42mmoles HNO3 You have the volume and you have the quantity of HNO3. V'=30.42 mL n'=42.5 mmoles cM'=n'/V'=30.42/42.5=0.71576 M